# Three-Dimensional to Layered Halide Perovskites: A Parameter-Free Hybrid Functional Method for Predicting Electronic Band Gaps

### v1

@@ Abstract
In order to quantify the performance of the different functionals in describing the band gap of layered halide perovskites, we calculate the mean absolute error (MAE) and root mean square error (RMSE). Figure 4a summarizes the performance of all hybrid functionals for all the layered materials, as well as by halogen (A2PbCl4, A2PbBr4, A2PbI4). The HSE functional is the least suitable, with errors close to 1.0 eV. PBE0, on the other hand, predicts the band gap reasonably well across the layered materials, particularly for iodides. The proposed DSH0() has a similar performance and gives the most minor errors overall for bromides and iodides. In particular, the MAE and RMSE are below 80 meV for iodine-based materials. We also include two Dion-Jacobson layered halide perovskites (see Table S3), further validating the observed performance trend for all layered materials. Since in DSH(), a fixed Hartree fraction is adopted to set the screening in layered perovskites with the same halogen. Consequently, any variation of the band gap of materials with the same inorganic layer can be attributed to structural distortions of the PbX6 octahedra due to the different organic cations. For instance, PEA2PbBr4 and PMA2PbBr4 have similar Pb-X-Pb bond angles of 151∘ and 150∘ respectively, resulting in practically identical band gaps of 3.50 eV. While BA2PbBr4, which exhibits slightly less octahedral distortion with the Pb-X-Pb angle amounting to 155∘, the band gap is narrower 3.26 eV as the orbital overlap increases. A similar trend also applies to the iodides. This relationship between the increase in Pb-X-Pb distortion with the increase in the band gap is well documented and understood both theoretically 3 and experimentally 62.
@@
Finally, since the DSH() is successfully predicting the band gaps of both 3D and the different types of layered halide perovskites, we extend our analysis to verify that this approach can also predict band gaps for layered materials with a thicker inorganic layer (i.e., larger -numbers). Thus, we apply our method to , where is the number of octahedrons in the inorganic layer. We find that the band gap decreases as increases, which is expected due to the reduction of the quantum, and dielectric confinement 65, 10, 3. We also note that to avoid introducing an artificial dipole due to the specific orientation of the organic molecule in the perovskite lattice, for and , we replace in with Cs (see the inset of Figure 4). The calculated band gaps employing PBE0 and DSH() are compared to experimental electronic band gaps shown in Figure 4. For the 3D case, we employ the DSH experimental band gaps of p-CsPbI3. As shown in Figure 4, the DSH() band gaps agree well with the experiments for the complete series of . Notably, our band gap for is comparable to the GW band gap reported by Giorgi et al. 16 and the experimental values reported in Table S4. This comparison also highlights the predictive accuracy of DSH(. It offers a useful framework that can be used for other layered halide perovskites and their heterostructures with other 3D materials.
@@
Overall, we present results for the band gaps of 3D and layered halide perovskite materials employing a consistent workflow to optimally set up the doubly screened dielectric-dependent hybrid (DSH) functional. Our results demonstrate that the proposed DSH() functional, which incorporates only material-dependent mixing parameters like the macroscopic high-frequency dielectric constants, predicts the band gaps with better accuracy than standard hybrid functionals in both 3D and layered perovskites. Our findings underline the importance of accounting for local disorder when benchmarking any functional for 3D halide perovskites. For the layered materials, it is evident that special attention needs to be taken when setting the screening based on the macroscopic dielectric function, for example, DSH0() and DSH(), using the average of the diagonal components of the dielectric constant matrix tend to overestimate the measure band gaps. However, we used the DSH functional with the dielectric constants of the corresponding 3D perovskites, DSH(), described at room-temperature with polymorphous structures, to predict band gaps. We achieve remarkable performance in predicting experimental electronic band gaps across all the layered perovskites considered, irrespective of the type of organic spacer, especially for iodine-based materials. We observe that for all the materials (i.e., 3D and layered), the HSE functional significantly and systematically underestimates the band gaps due to the absence of non-local long-range dielectric screening. On the other hand, the PBE0 functional that shares the same non-local short-range screening with HSE but has an additional 25% non-local long-range screening performs very well. It could be the functional of choice if one cannot afford to compute the dielectric properties of halide perovskites. The computational strategy presented here sets up a parameter-free ab initio methodology that allows efficient prediction of the electronic properties of both 3D and layered halide perovskites. As such, the proposed workflow and benchmark can be extremely powerful in investigating 3D/2D heterostructures, which are challenging to describe at the ab initio level, while being attractive to develop stable and efficient opto-electronic devices.
@@
I.B.G acknowledges discussionsns with Cui Zhi-Hao on the implementation of DSH functional in the VASP code. G.V. and I.B.G. acknowledge funding from the Agence Nationale pour la Recherche through the CPJ program and the SURFIN project (ANR-23-CE09-0001). This work was granted access to the HPC resources of TGCC under the Allocation Grant No. 2022 - A0130907682 made by GENCI and Lumi. M.Z. was funded by the European Union (project ULTRA-2DPK / HORIZON-MSCA-2022-PF-01 / Grant Agreement No. 101106654). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them. We acknowledge computational resources from the EuroHPC Joint Undertaking and supercomputer LUMI [https://lumi-supercomputer.eu/], hosted by CSC (Finland) and the LUMI consortium through a EuroHPC Extreme Scale Access call.
@@
Doubly screened dielectric dependent hybrid functional; Computational details; Comparison of model dielectric functions in hybrid functionals; Lattice constants and band gaps of 3D perovskites calculated using different functionals in Figure 2; Band gaps of layered halide perovskites calculated using different functionals in Figure 3; Band gaps of BA series in Figure 4; Comparison of model dielectric functions for hybrid functionals in 3D perovskites; Comparison of model dielectric functions for hybrid functionals in layered halide perovskites; Convergence of dielectric constant and band gap with self-consistent DSH iteration in 3D perovskites

### v2

@@ Abstract
The proposed DSH() functional achieves the highest accuracy, yielding the lowest MAE (RMSE) of 0.14 eV (0.18 eV). For all iodides, PBE0 exhibits the same accuracy, which can be practical for future calculations as one does not need to first calculate . However, for perovskites with higher band gaps (i.e., bromides and chlorides) DSH outperforms in terms of accuracy. Overall, PBE0, DSH0(), DSH(), and HSE, exhibit MAE (RMSE) values of 0.22 eV (0.26 eV), 0.40 eV (0.45 eV), 0.52 eV (0.56 eV), and 0.84 eV (0.85 eV), respectively. We also include two Dion-Jacobson layered halide perovskites (see Table S3), further validating the observed performance trend for all layered materials. Since in DSH(), a fixed Hartree fraction is adopted to set the screening in layered perovskites with the same halogen. Consequently, any variation of the band gap of materials with the same inorganic layer can be attributed to structural distortions of the PbX6 octahedra due to the different organic cations. For instance, PEA2PbBr4 and PMA2PbBr4 have similar Pb-X-Pb bond angles of 151∘ and 150∘ respectively, resulting in practically identical band gaps of 3.50 eV. While BA2PbBr4, which exhibits slightly less octahedral distortion with the Pb-X-Pb angle amounting to 155∘, the band gap is narrower 3.26 eV as the orbital overlap increases. A similar trend also applies to the iodides. This relationship between the increase in Pb-X-Pb distortion with the increase in the band gap is well documented and understood both theoretically [27] and experimentally [28].
@@
Finally, since the DSH() is successfully predicting the band gaps of both 3D and the different types of layered halide perovskites, we extend our analysis to verify that this approach can also predict band gaps for layered materials with a thicker inorganic layer (i.e., larger -numbers). Thus, we apply our method to , where is the number of octahedrons in the inorganic layer. We find that the band gap decreases as increases, which is expected due to the reduction of the quantum, and dielectric confinement [52, 6, 27]. We also note that to avoid introducing an artificial dipole due to the specific orientation of the organic molecule in the perovskite lattice, for and , we replace in with Cs (see the inset of Figure 4). The calculated band gaps employing PBE0 and DSH() are compared to experimental electronic band gaps shown in Figure 4. For the 3D case, we employ the DSH experimental band gaps of p-CsPbI3. As shown in Figure 4, the DSH() band gaps agree well with the experiments for the complete series of . Figure S5 shows the DOS for all BA compounds. Similar to the other layered materials, the band edges are made of orbital contributions from the inorganic framework. Notably, our band gap for is comparable to the GW band gap reported by Giorgi et al. [19] and the experimental values reported in Table S4. This comparison also highlights the predictive accuracy of DSH(. It offers a useful framework that can be used for other layered halide perovskites and their heterostructures with other 3D materials.
@@
Overall, we present results for the band gaps of 3D and layered halide perovskite materials employing a consistent workflow to optimally set up the doubly screened dielectric-dependent hybrid (DSH) functional. Our results demonstrate that the proposed DSH functional, which incorporates only material-dependent mixing parameters based on macroscopic high-frequency dielectric constants, predicts band gaps more accurately than standard hybrid functionals for both 3D and layered perovskites. Our findings highlight the importance of accounting for local disorder when benchmarking any functional for 3D halide perovskites. For layered materials, it is evident that special attention must be given when setting the screening based on the dielectric constant since using the average of the diagonal components of the dielectric constant matrix in DSH0() and DSH(), tends to overestimate the measured band gaps. Consequently, employing the DSH functional with dielectric constants from locally disordered 3D perovskite structures DSH() to predict band gaps gives the highest accuracy with MAE of 0.14 eV across all layered perovskites, regardless of the organic spacer type, or the halogen in the inorganic framework. In contrast, the HSE functional significantly underestimates band gaps due to the lack of non-local long-range dielectric screening, resulting in the highest MAE of 0.84 eV for all materials. On the other hand, the PBE0 functional that shares the same non-local short-range screening with HSE but has an additional 25% non-local long-range screening performs very well, with MAE of 0.22 eV. It could be the functional of choice if one cannot afford to compute the dielectric properties of halide perovskites. The computational strategy presented here sets up a parameter-free ab initio methodology that allows efficient prediction of the electronic properties of both 3D and layered halide perovskites. As such, the proposed workflow and benchmark can be extremely powerful in investigating 3D/2D heterostructures, which are challenging to describe at the ab initio level, while being attractive to develop stable and efficient opto-electronic devices.
Acknowledgment
@@
I.B.G acknowledges discussions with Cui Zhi-Hao on the implementation of the DSH functional. We acknowledge funding from the Agence Nationale pour la Recherche through the CPJ program and the SURFIN project (ANR-23-CE09-0001), the ALSATIAN project (ANR-23-CE50-0030), and the ANR under the France 2030 programme, MINOTAURE project (ANR-22-PETA-0015). M.Z. was funded by the European Union (project ULTRA-2DPK / HORIZON-MSCA-2022-PF-01 / Grant Agreement No. 101106654). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them. We acknowledge computational resources from the EuroHPC Joint Undertaking and supercomputer LUMI [https://lumi-supercomputer.eu/], hosted by CSC (Finland) and the LUMI consortium through a EuroHPC Extreme Scale Access call, and access to the HPC resources of TGCC under the Allocation Grant No. 2022 - A0130907682 made by GENCI
Supporting Information Available
@@
Doubly screened dielectric dependent hybrid functional; Computational details; Comparison of model dielectric functions in hybrid functionals; Lattice constants and band gaps of 3D perovskites calculated using different functionals in Figure 2; Band gaps of layered halide perovskites calculated using different functionals in Figure 3; Band gaps of BA series in Figure 4; Comparison of model dielectric functions for hybrid functionals in 3D perovskites; Comparison of model dielectric functions for hybrid functionals in layered halide perovskites; Convergence of dielectric constant and band gap with self-consistent DSH iteration in 3D perovskites; Density of states of layered halide perovskites and BA2MAn-1PbnI3n+1 ( n = 1, 2, 3).